OpenStudy (anonymous):

the number of traffic accidents that occurs on a particular road during a month follows a poison distribution with mean of 3.what is the probability that less than 2months until next traffic accidents is occured?

OpenStudy (kropot72):

\[\large P(X=x)=\frac{e^{-\lambda} \lambda^{x}}{x!}\] \[\large P(X<2)=P(X=0)+P(X=1)\] \[\large P(X<2)=\frac{e^{-3} 3^{0}}{0!}+\frac{e^{-3}3^{1}}{1!}=you\ can\ calculate\]